### v2

@@ 4.2 Embedding Reservation for Few-Shot Adaptation
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as [47]. $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}\circ h_{1}(\mathbf{x})$.
Third, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
$\mathcal{L}_{\mathrm{v}}=$ $F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{ce}(f_{0}(\mathbf{z}),\hat{y})+$
$+$ $F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}}).$ (4)
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{p}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss [46], and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
@@
where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector. The final loss for training the initial defense model is:
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\quad\mathcal{L}_{0}=\mathcal{L}_{\mathrm{ce}}+\gamma\mathcal{L}_{\mathrm{virtual}}$ (6)
In Sec. 4.2, it forces an instance to be away from the reserved virtual classes (item 1) and pushes the mixed instance towards a virtual class (item 2) and away from other classes (item 3). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved [47]. The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot feedback.
4.3 Prototype Augmentation for Memory Efficiency
@@
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [50]. Many CL methods store a fraction of old data to jointly train the model with current data to address this issue [27]. However, preserving old feedback may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation [51] to maintain the decision boundary of previous stages, without saving any feedback. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
@@
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ the same as ProtoNet [30]:
@@
where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of feedback for each class.
In each stage, the defense model is adapted using the following loss:
@@
4.4 Model Ensemble for Clean and Adversarial Image Classification
To maintain high classification accuracy on both clean and adversarial images mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) [34] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean images. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models
We adopt the self-perturbation [37] to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. Details of the self-perturbation are given in the supplementary.
The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
@@ 4.4 Model Ensemble for Clean and Adversarial Image Classification
The overall algorithm of CAD framework is presented in Algorithm 1.
@@
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.

### v3

@@ Introduction
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as (Zhou et al. 2022). $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}\circ h_{1}(\mathbf{x})$.
Finally, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
$\mathcal{L}_{0}=$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})+$
$+$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}})+F_{\mathrm{ce}}(f_{0}(\mathbf{z}),\hat{y}).$ (4)
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{p}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss (Zhang and Sabuncu 2018), and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
@@
where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector.
In Eq. Embedding Reservation for Few-Shot Adaptation, it pushes an instance away from the reserved virtual classes (item 1) while towards known classes (item 2), and forces a virtual instance away from known class (item 3) while towards virtual class (item 4). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved (Zhou et al. 2022). The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot budget.
Prototype Augmentation for Memory Efficiency
@@
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased (Zhu et al. 2021). Many CL methods store a fraction of old data to jointly train the model with current data to address this issue (Rebuffi et al. 2017). However, preserving old budget may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation (Zhu et al. 2022) to maintain the decision boundary of previous stages, without saving any budget. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
@@
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ following ProtoNet (Snell, Swersky, and Zemel 2017):
@@
where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of budget for each class.
At each stage, the defense model is adapted using the following loss:
@@
Model Ensemble for Clean Data Classification
To maintain high classification accuracy on both clean and adversarial data mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) (Tramèr et al. 2018) which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean data. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models.
We adopt self-perturbation (Wang et al. 2023a) to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
@@
In this way, both clean and adversarial images are correctly classified.
For the overall algorithm of the CAD framework, details about self-perturbation, and limitations of CAD, please refer to the supplementary materials.
Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies comparing CAD with baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy (Acc) against each attack after adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy (AIAcc) (Rebuffi et al. 2017), for assessing the knowledge retention of the model.
Datasets
Experiments are conducted on CIFAR-10 and ImageNet-100. CIFAR-10 (Krizhevsky, Hinton et al. 2009) contains 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 (Tian, Krishnan, and Isola 2020) is a subset of ImageNet (Deng et al. 2009) and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attack Algorithms
We choose PGD-$l_{\infty}$ (Madry et al. 2017) as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM (Kurakin, Goodfellow, and Bengio 2016), RFGSM (Tramèr et al. 2018), MIM (Dong et al. 2018), DIM (Wu et al. 2021), NIM (Lin et al. 2019), SNIM (Lin et al. 2019), VNIM and VMIM (Wang and He 2021) to compose the attack pool in defense performance evaluation. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al. 2019). As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM and SNIM, $l_{2}$ attack CW (Carlini and Wagner 2017) and DeepFool (Moosavi-Dezfooli, Fawzi, and Frossard 2016), $l_{1}$ attack EAD (Chen et al. 2018) and EADEN (Chen et al. 2018), and $l_{0}$ attack OnePixel (Su, Vargas, and Sakurai 2019) and SparseFool (Modas, Moosavi-Dezfooli, and Frossard 2019) compose the attack pool. Details of attack parameters are given in the supplementary.
Stage:Attack 0:PGD 1:SNIM 2:BIM 3:RFGSM 4:MIM 5:DIM 6:NIM 7:VNIM 8:VMIM Clean
None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
TRADES (Zhang et al. 2019) 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
GAIRAT (Zhang et al. 2020) 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
RPF (Dong and Xu 2023) 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
DMAT (Wang et al. 2023b) 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
ADP (Yoon, Hwang, and Lee 2021) 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure (Nie et al. 2022) 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
Table 1: Classification accuracy(%) against attacks($l_{\infty}$) on ImageNet-100. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD in case no constraints on memory and defense budget, and is considered to be the empirical upper bound marked with a gray background.
Stage:Attack 0:PGD 2:SNIM 3:VNIM 4:CW 5:DeepF. 6:EAD 7:EADEN 8:OnePixel 9:SparseF.
None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
EWC (Kirkpatrick et al. 2017) 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE (Zhu et al. 2022) 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF (Wang et al. 2022) 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT (Zhou et al. 2022) 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
Table 2: Continual defense performance (average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10. DeepF. and SparseF. denote DeepFool and SparseFool respectively.
Implementation Details
Both the target model and defense model use ResNet (He et al. 2016) as the backbone network in our method. Specifically, we use WideResNet-28-10 for CIFAR-10 following (Grathwohl et al. 2020) and ResNet-50 for ImageNet-100 following (Laidlaw, Singla, and Feizi 2020). A 4-layer ConvNet is used as the weight estimator model. The cosine classifier (Gidaris and Komodakis 2018) is adopted as the classifier of the defense model. We utilize Torchattacks (Kim 2020) to generate adversarial images. Shots of defense budget per class is set to $K=10$. The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following (Zhou et al. 2022). Epochs for training and fine-tuning are set to $100$ and $4$ respectively.
Defense Baselines
We first compare with adversarial training methods TRADES (Zhang et al. 2019), GAIRAT (Zhang et al. 2020), RPF (Dong and Xu 2023) and DMAT (Wang et al. 2023b), and purification methods ADP (Yoon, Hwang, and Lee 2021) and DiffPure (Nie et al. 2022) in defense performance evaluation. As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC (Kirkpatrick et al. 2017), SSRE (Zhu et al. 2022), BEEF (Wang et al. 2022), and FACT (Zhou et al. 2022). Besides, we propose another variant of our method as the empirical upper bound, denoted as premium-CAD. For premium-CAD, there is no constraint of memory and no concerns about the amount of defense budget. We train additional attack-specific defense models using abundant defense budget in each stage, and ensemble them with the target model by the weight estimator. For details about the implementation of defense baseline and premium-CAD, please refer to supplementary materials.
Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. As depicted in Table 1, despite the small scale of the defense budget, CAD demonstrates strong adaptability to each type of attack. When compared with the top baseline DMAT and ADP on ImageNet-100, CAD outperforms them by approximately $6\%$ for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $88.1\%$ accuracy, while the target model stands at $89.8\%$. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
(a) PGD
(b) CW
(c) EAD
(d) OnePixel
Figure 3: T-SNE diagrams of features of test data encoded by the defense model after all stages. Each color represents a class of CIFAR10. After several steps of adaptation, CAD retains the capability to categorize the attacked data.
To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods on CIFAR-10 in a dynamic scenario formed by attacks under different norms $l_{p=0,1,2,\infty}$. As depicted in Table 2, CAD maintains robust defense performance despite changes in attack types. However, as the variety of attack types increases, the defense performance gradually diminishes. This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense. Compared with the top baseline method FACT, CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense budget and memory resources, premium-CAD emerges as the superior choice.
Ablation Study
As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting. Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images.
Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel Clean
w/o $\mathcal{L}_{0}$ 94.5 82.4 76.1 71.8 95.8
w/o $\mathcal{L}_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
CAD(ours) 94.5 89.9 86.6 85.3 95.8
Table 3: AIAcc(%) on CIFAR-10 for ablation.
Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and OnePixel($l_{0}$) to validate the efficacy of each component. As shown in Table 3, CAD without $\mathcal{L}_{0}$ and $\mathcal{L}_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage. The absence of model ensemble leads to the inability to uphold performance on clean images.
Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Figure 3 to visualize the defense capability of CAD after adaptation. These features are encoded by the defense model after the last stage of adapting to OnePixel attack. From (a) to (d) in Figure 3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types. After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
Multiple Attacks in One Stage
Stage 0 1 2 3
$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
Table 4: Acc(%) against multiple attacks in one stage.
In practice, it is common for multiple attacks to occur simultaneously within a single stage. To assess CAD’s adaptability in such scenarios, on CIFAR-10, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks per stage. As presented in Table 4, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
Number of Shots for Adaptation
Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel
$K=15$ 94.5 90.0 88.4 85.5
$K=10$ 94.5 89.9 86.6 85.3
$K=5$ 94.5 89.4 86.1 84.6
$K=1$ 94.5 87.9 84.6 83.4
Table 5: AIAcc(%) with different budget scale.
As indicated in Principle 2, the defense adaptation relies on few-shot defense budget. To investigate the influence of the budget shots $K$, we conducted tests with $K=15,10,5,1$ on CIFAR-10. The results presented in Table 5 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($83.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
Memory Allowance for Adaptation
Cache Size Model Size
Stage 1 2 3 1 2 3
CAD 50.3K 75.5K 100.4K 141M 141M 141M
$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
stage on CIFAR-10. As depicted in Table 6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
Conclusion
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect few-shot defense budget from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles that include continual adaptation without catastrophic forgetting, few-shot budget adaptation, memory-efficient adaptation, and high accuracy for both clean and adversarial data, and CAD demonstrates strong performance when operating under these principles. Through extensive experiments, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
Appendix
@@ Conclusion
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot budget $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.